What is the final temperature when 100 g of iron at 90 degrees Celsius is placed in 400 g of water at 25 degrees Celsius? The specific heat capacity of iron 0.45 J/g*K and the specific heat capacity of water is 4.184 J/g*K.
Answer #1

let final temp be T
heat gained by water = mCdT = 400*4.184*(T - 25)
heat glost by iron = 100*0.45*(90 - T)
Heat lost = heat gained [conservation of energy]

400*4.184*(T-25) = 100*0.45*(90 - T)
37.1911T - 929.77 = 90 - T
T = 26.701 degrees